A gold smith combined an alloy that costs $430 per gram with an alloy that costs $180 per gram. How many grams of each were used to make a mixture of 200 grams that cost $250 per gram?
Answer by stanbon(75887)

A gold smith combined an alloy that costs $430 per gram with an alloy that costs $180 per gram. How many grams of each were used to make a mixture of 200 grams that cost $250 per gram?
---
value + value = value
430x + 180(200-x) = 250*200
430x +180*200 - 180x = 250*200
250x = 70*200
x = 56 grams (amt. of $430 alloy in the mixture)
200-56 = 144 grams (amt. of $180 alloy in the mixture)
========================================================
Cheers,
Stan H.